You were hired as a consultant to Quigley Company, whose target capital structure is 40% debt, 10% preferred, and 50% common equity. The interest rate on new debt is 6.50%, the yield on the preferred is 7.50%, the cost of retained earnings is 13.25%, and the tax rate is 34%. The firm will not be issuing any new stock. What is Quigley's WACC? Enter your answer rounded to two decimal places. Do not enter % in the answer box. For example, if your answer is 0.12345 or 12.345% then enter as 12.35 in the answer box.
Answer #1

Weight of capital

Debt = 40%

Preferred stock = 10%

Common equity = 50%

Cost of debt Kd = 6.5%

cost of equity = cost of retained earning Ke = 13.25%

Cost of preferred stock Kp = 7.5%

tax rate = 34%

using WACC = Wd*Kd*(1-t) + We*Ke + Wp*Kp = 0.4*6.5*(1-0.34) + 0.5*13.25 + 0.1*7.5 = 9.09%

Quigley's WACC is 9.09%